OpenStudy (anonymous):

Given the function g(x) = 3log3^x What is/are the zero(s) of the function?

OpenStudy (anonymous):

3log3^x=0 log3^x=0 3^x=2^0 a^0=1 any thing^0=1 so 1^0=1 2^0=1 3^0=1 3^x=1 3^x=3^0 nw find x

OpenStudy (anonymous):

so the zero is 1

OpenStudy (anonymous):

x=0 g(x)=3log1 =0